# Continual Vision-and-Language Navigation

### v2

@@ Abstract
In developing Vision-and-Language Navigation (VLN) agents that navigate to a destination using natural language instructions and visual cues, current studies largely assume a train-once-deploy-once strategy. We argue that this kind of strategy is less realistic, as deployed VLN agents are expected to encounter novel environments continuously through their lifetime. To facilitate more realistic setting for VLN agents, we propose Continual Vision-and-Language Navigation (CVLN) paradigm for agents to continually learn and adapt to changing environments. In CVLN, the agents are trained and evaluated incrementally across multiple scene domains (i.e., environments). We present two CVLN learning setups to consider diverse forms of natural language instructions: Initial-instruction based CVLN, focused on navigation via initial-instruction interpretation, and dialogue-based CVLN, designed for navigation through dialogue with other agents. We introduce two simple yet effective baseline methods, tailored to the sequential decision-making needs of CVLN: Perplexity Replay (PerpR) and Episodic Self-Replay (ESR), both employing a rehearsal mechanism. PerpR selects replay episodes based on episode difficulty, while ESR stores and revisits action logits from individual episode steps during training to refine learning. Experimental results indicate that while existing continual learning methods are insufficient for CVLN, PerpR and ESR outperform the comparison methods by effectively utilizing replay memory.
@@ 1 Introduction
One of the desired attributes of robotic agents is their ability to learn and adapt to new environments without losing knowledge of previous ones. In practice, robots deployed in the human world are required to perform a variety of tasks in complex and changing environments. Vision-and-Language Navigation (VLN) focuses on agents that navigate using natural language instructions and visual cues, promising for applications like robotic assistants. Despite recent advancements, most VLN research follows a train-once-deploy-once strategy.
In real-world applications, environments are not static; they evolve over time with new buildings, renovations, or changes in layout. A deployed VLN agent must promptly adapt whenever the environment changes throughout its lifetime. Finetuning on new environments might seem practical, but it can cause the agent’s performance in previously learned contexts to deteriorate, a challenge known as catastrophic forgetting [30]. To address this challenge, the agent needs the capability for Continual Learning (CL) [42], which enables it to learn and adapt to new circumstances while retaining its existing knowledge.
In this paper, we propose Continual Vision-and-Language Navigation (CVLN) paradigm, which trains and evaluates VLN agents with consideration of the CL ability. In CVLN, agents explore different scene domains—each comprising various indoor scenes with multiple navigation episodes—sequentially and are evaluated on their ability to navigate through all of them. The CVLN agents need to retain their ability on previously learned scene domains while learning the current scene domain. Figure 2 illustrates an overview of CVLN. In CVLN, the agent incrementally learns multiple scene domains, and as shown in Figure 1, there are substantial differences in visual distributions among scene domains. These distributional discrepancies can lead to catastrophic forgetting when the domains are learned incrementally.
@@
The CVLN supports two settings to handle different types of natural language instructions: Initial-instruction based CVLN (I-CVLN) and Dialogue based CVLN (D-CVLN). In I-CVLN, the instructions given to the agent at the start of navigation contain fine-grained information about the entire navigation path. In D-CVLN, the agent receives a target object and a conversation history between humans collaborating to locate it, and must reason navigation actions to reach the goal within environments.
@@
We present two simple yet effective baseline methods for CVLN—Perplexity Replay (PerpR) and Episodic Self-Replay (ESR)—both based on a rehearsal mechanism, where the effective construction and utilization of replay memory are crucial. PerpR uses self-evaluated action perplexity (i.e. difficulty) to assess the importance of episodes, effectively selecting and organizing them in the replay memory based on their action perplexity. After learning a scene domain, the agent adds high-perplexity episodes to the replay memory and removes low-perplexity ones. Meanwhile ESR stores action logits predicted by the agent for each individual step within episodes in the replay memory. During subsequent training, the agent revisits these stored logits, allowing it to refine its learning by leveraging detailed episode step-level logits from past experiences.
To validate the effectiveness of our methods, we conducted comparative experiments with existing CL methodologies. Furthermore, we conducted an ablation study, analyzed the effects of varying replay memory sizes, and examined the stability-plasticity trade-off.
Figure 1: Frechet Inception Distance[16] heatmaps illustrating the distribution shifts between visual features [47] across 10 scene domains for the I-CVLN and D-CVLN.
To summarize our contributions:
  • We propose a Continual Vision-and-Language Navigation (CVLN) paradigm to develop VLN agents capable of continual learning, introducing two settings— I-CVLN (instruction-based) and D-CVLN (dialogue-based)—for comprehensive evaluation across diverse natural language instruction types.
  • We propose two simple yet effective CVLN methods, PerpR and ESR, which enhance continual learning through rehearsal mechanisms, applicable to tasks like CVLN that involve sequential decision-making.
  • In experiments, PerpR and ESR demonstrate their effectiveness in the CVLN paradigm, achieving improved performance over comparative methods in both I-CVLN and D-CVLN settings.
Figure 2: Comparison between (a) VLN and (b) CVLN. In VLN, the agent is trained within fixed scene domains and then evaluated on unseen scene domains. In contrast, the agent in CVLN is trained within a new scene domain sequentially and and evaluated on those scene domains along with previously encountered ones.
@@ 2 Related Work
Continual Learning. Continual Learning (CL) focuses on training models to acquire new knowledge over time while retaining previously learned information. CL methods have generally been extensively studied for simple tasks [13] in computer vision. This field includes several CL scenarios. Instance-Incremental Learning trains on batches of samples for the same task [27]. Domain-Incremental Learning (DIL) deals with tasks having the same labels but different input distributions, with no domain identity during inference [19]. Task-Incremental and Class-Incremental Learning involve different label spaces, with varying use of domain identity [19]. Task-Free Continual Learning handles separate label spaces without domain identity [1], and Online Continual Learning processes sequential data streams in real-time [2]. CVLN fits within DIL, providing scene domain identities during training but not during testing, requiring generalization without explicit domain knowledge at evaluation. There are three main approaches for CL: Regularization [22], which adds terms to prevent the loss of prior knowledge; Rehearsal [34, 9, 8, 7], which stores and reuses data from past tasks; and Architectural [36, 29], which introduces task-specific parameters to the agent’s architecture. This paper proposes two Rehearsal-based methods, PerpR and ESR, and evaluates their effectiveness through comparative experiments.
@@
Vision-and-Language Navigation. In VLN, agents navigate realistic 3D indoor environments by following natural language instructions. Various datasets like R2R [6], R4R [20], and RxR [24] have been developed, focusing on navigation based on initial instructions that fully describe the path. Other tasks, such as CVDN [41] and HANNA [32], involve agents obtaining navigation information through dialogue with an oracle or other agents. Additionally, datasets like Touchdown [10] and StreetLearn [31] have expanded VLN research to outdoor environments. IVLN [23] present a paradigm where agents maintain memory across episodes to navigate a sequence of episodes. In the field of VLN, research focuses on improving agent architecture [17, 11, 14, 4], enhancing visual-text alignment through auxiliary losses [28, 46, 48], using data augmentation [40, 45, 25, 21], and extensive pre-training [15, 11, 12, 3] to enhance generalization. In this paper, we adopted VLN-BERT [17] and HAMT [11] as the backbone architectures. Existing VLN datasets and methods lack considerations for CL, highlighting the need for scalable strategies that enable agents to learn efficiently to diverse, changing environments without costly retraining models from scratch.
@@ 3 Continual Vision-and-Lauguage Navigation
3.1 Formulation of VLN
The existing VLN agents are trained on fixed training data and evaluated on episodes from scenes that were not encountered during training. The objective of VLN agents is to minimize the following loss:
$\mathcal{L}_{D}\triangleq\mathbb{E}_{(I,A^{*})\sim D}[\ell(\pi_{\theta}(V,I),A^{*})]$ (1)
where $\theta$ specifies the parameters for the VLN agent, $\pi$. The dataset $D$ is composed of episodes, which include navigation instructions $I$ and their corresponding ground-truth trajectory $A^{*}=\{a_{1}^{*},a_{2}^{*},\ldots,a_{N}^{*}\}$. The set $V=\{v_{1},v_{2},\ldots,v_{N}\}$ represents the sequence of visual observations from the environment, each associated with a step determined by a predicted action, with $N$ denoting the total number of steps per episode. The agent $\pi_{\theta}(V,I)$ determines the actions based on $V$ and $I$ sequentially. The navigation imitation loss function $\ell$ then compares this behavior to the ground-truth trajectory $A^{*}$ to evaluate the discrepancy between them. We use cross entropy loss for $\ell$.
3.2 Formulation of CVLN
A scene domain includes episodes in multiple indoor environments, and each episode comprises a natural language navigation instruction and its corresponding trajectory. CVLN agent learns scene domains incrementally. The agent is evaluated on the incrementally learned scene domains. The evaluation episodes belong to the learned scene domains but are episodes that have not been seen during training. During this process, the agent must learn the current scene domain without forgetting the knowledge about previously learned scene domains. This learning approach is more suitable for real-world agents that encounter and need to learn new scene domains continually.
Formally, a CVLN is divided into $S$ scene domains; during each scene domain $s\in\{1,\ldots,S\}$, an episode $\epsilon$, navigation instruction $I$ and their corresponding ground-truth trajectory $A^{*}$ are drawn from an independent and identically distributed (i.i.d.) distribution $D_{s}$. The objective of agents in CVLN is as follows:
@@ 3.2 Formulation of CVLN
$\quad\text{where}\ \mathcal{L}_{D_{s}}\triangleq\mathbb{E}_{(I,A^{*})\sim D_{s}}[\ell(\pi_{\theta}(V,I),A^{*})]$
@@
3.3 Initial-Instruction based CVLN
In I-CVLN, agents receive initial instructions outlining the entire navigation path. As the R2R [6] and RxR [24] datasets provide initial instruction-based navigation episodes, we construct our continual learning benchmark for I-CVLN on top of these datasets, using only english epidodes. The training split from the original datasets is used for I-CVLN training, while the val-seen split, which contain episodes from the same scenes but not seen during training, is used for I-CVLN validation. We grouped episodes from the same scenes across both datasets into 20 scene domains, each containing over 600 episodes for training.
Datasets Split Episode Scene Domain
I-CVLN Train 15700 20
Valid 1563 20
D-CVLN Train 3737 11
Valid 382 11
Table 1: Dataset statistics for I-CVLN and D-CVLN. The table outlines the number of episodes and scene domains available in the training (Train) and validation (Valid) splits for each dataset.
3.4 Dialogue based CVLN
In D-CVLN, agents navigate from dialogue history. As the CVDN [41] dataset includes dialogue data between a follower and an oracle paired with corresponding trajectories, we build benchmark for D-CVLN on top of it. Due to the limited episodes per scene in CVDN, we grouped four scenes into each scene domain, resulting in 11 domains for training. These scene domains are learned sequentially. The training split of CVDN is used for D-CVLN training, while the val-seen split, containing evaluation episodes for each domain but not seen during training, is used for D-CVLN validation. Figure 3 illustrates a comparison between I-CVLN and D-CVLN.
3.5 Evaluation Protocol for CVLN
The evaluation of CVLN agents focuses on assessing their ability to retain previously learned knowledge while effectively adapting to new scene domain. For CVLN evaluation, the Average Metric is used to evaluate agent performance across different scene domains. The average metric is defined as follows:
@@ 3.5 Evaluation Protocol for CVLN
where $R_{S,i}$ indicates the agent’s performance on the $i$th domain after training the $S$th scene domain. The metric evaluates the agent’s knowledge retention across prior scene domains and its adaptability to new ones.
For I-CVLN, we use standard VLN metrics [5] $R_{S,i}$ in Equation 3: Success Rate (SR), Success weighted by inverse Path Length (SPL), and Navigation Error (NE). For D-CVLN, we use the main metric from CVDN [41], $R_{S,i}$ in Equation 3, which is Goal Progress (GP) in meters.
@@ 4 Methods
Rehearsal-based CL approaches mitigate catastrophic forgetting by using a fixed-size replay memory $M$ to store and revisit episodes from previous domains. After training each task, $M$ is updated by deleting an equal number of old episodes and adding new ones from the latest domain. Effective replay memory construction and usage are crucial for maintaining performance across domains. In this context, we propose Perplexity Replay (PerpR) and Episodic Self-Replay (ESR) as baseline methods for the CVLN agent. Figure 4 provides an overview of PerpR and ESR.
Figure 4: Overview of Perplexity Replay (PerpR) and Episodic Self-Replay (ESR) for CVLN agents. (a) PerpR prioritizes challenging episodes in the agent’s memory, optimizing for high Action Perplexity (AP). (b) ESR enables the agent to self-replay using past optimal behaviors.

### v3

@@ Abstract
Developing Vision-and-Language Navigation (VLN) agents typically assumes a train-once-deploy-once strategy, which is unrealistic as deployed agents continually encounter novel environments. To address this, we propose the Continual Vision-and-Language Navigation (CVLN) paradigm, where agents learn and adapt incrementally across multiple scene domains. CVLN includes two setups: Initial-instruction based CVLN for instruction-following, and Dialogue-based CVLN for dialogue-guided navigation. We also introduce two simple yet effective baselines for sequential decision-making: Perplexity Replay (PerpR), which replays difficult episodes, and Episodic Self-Replay (ESR), which stores and revisits action logits during training. Experiments show that existing continual learning methods fall short for CVLN, while PerpR and ESR achieve better performance by efficiently utilizing replay memory.
@@ 1 Introduction
One of the desired attributes of robotic agents is their ability to learn and adapt to new environments without losing knowledge of previous ones. In practice, robots deployed in the real world continually encounter unseen environments and are required to perform a variety of tasks in them. Vision-and-Language Navigation (VLN) focuses on agents that navigate using natural language instructions and visual cues, promising for applications like robotic assistants. Despite recent advancements, most VLN research follows a train-once-deploy-once strategy. A deployed VLN agent must promptly adapt whenever the environment changes throughout its lifetime. Finetuning on new environments might seem practical, but it can cause the agent’s performance in previously learned contexts to deteriorate, a challenge known as catastrophic forgetting [McCloskey and Cohen(1989)]. To address this challenge, the agent needs the capability for Continual Learning (CL) [Thrun and Pratt(2012)], which enables it to learn and adapt to new circumstances while retaining its existing knowledge.
@@
In this paper, we propose Continual Vision-and-Language Navigation (CVLN) paradigm, which trains and evaluates VLN agents with consideration of the CL ability. In CVLN, agents explore different scene domains—each comprising various indoor scenes with multiple navigation episodes—sequentially and are evaluated on their ability to navigate through all of them. The CVLN agents need to retain their ability on previously learned scene domains while learning the current scene domain. Figure 1 illustrates an overview of CVLN. The CVLN supports two settings to handle different types of natural language instructions: Initial-instruction based CVLN (I-CVLN) and Dialogue based CVLN (D-CVLN). In I-CVLN, the instructions given to the agent at the start of navigation contain fine-grained information about the entire navigation path. In D-CVLN, the agent receives a target object and a conversation history between humans collaborating to locate it, and must reason navigation actions to reach the goal within environments.
@@
We present two simple yet effective baseline methods for CVLN—Perplexity Replay (PerpR) and Episodic Self-Replay (ESR)—both based on a rehearsal mechanism, where the effective construction and utilization of replay memory are crucial. PerpR uses self-evaluated action perplexity (i.e. difficulty) to assess the importance of episodes, effectively selecting and organizing them in the replay memory based on their action perplexity. After learning a scene domain, the agent adds high-perplexity episodes to the replay memory and removes low-perplexity ones. Meanwhile ESR stores action logits predicted by the agent for each individual step within episodes in the replay memory. During subsequent training, the agent revisits these stored logits, allowing it to refine its learning by leveraging detailed episode step-level logits from past experiences. To validate the effectiveness of our methods, we conducted comparative experiments with existing CL methodologies. Furthermore, we conducted an ablation study, analyzed the effects of varying replay memory sizes, and examined the stability-plasticity trade-off.
Figure 1: Comparison between (a) VLN and (b) CVLN: VLN trains agents on fixed environments and evaluates them on unseen ones, while CVLN trains agents sequentially on new environments and evaluates them on both previously encountered and newly learned environments.
@@ 2 Related Work
Continual Learning. Continual Learning (CL) focuses on training models to acquire new knowledge over time while retaining previously learned information. CL methods have generally been extensively studied for simple tasks [Deng et al.(2009)Deng, Dong, Socher, Li, Li, and Fei-Fei] in computer vision. This field includes several CL scenarios. Instance-Incremental Learning trains on batches of samples for the same task [Lomonaco and Maltoni(2017)]. Domain-Incremental Learning (DIL) deals with tasks having the same labels but different input distributions, with no domain identity during inference [Hsu et al.(2018)Hsu, Liu, Ramasamy, and Kira]. Task-Incremental and Class-Incremental Learning involve different label spaces, with varying use of domain identity [Hsu et al.(2018)Hsu, Liu, Ramasamy, and Kira]. Task-Free Continual Learning handles separate label spaces without domain identity [Aljundi et al.(2019a)Aljundi, Kelchtermans, and Tuytelaars], and Online Continual Learning processes sequential data streams in real-time [Aljundi et al.(2019b)Aljundi, Lin, Goujaud, and Bengio]. CVLN fits within DIL, providing scene domain identities during training but not during testing, requiring generalization without explicit domain knowledge at evaluation. There are three main approaches for CL: Regularization [Kirkpatrick et al.(2017)Kirkpatrick, Pascanu, Rabinowitz, Veness, Desjardins, Rusu, Milan, Quan, Ramalho, Grabska-Barwinska, et al.], which adds terms to prevent the loss of prior knowledge; Rehearsal [Rebuffi et al.(2017)Rebuffi, Kolesnikov, Sperl, and Lampert, Chaudhry et al.(2018)Chaudhry, Ranzato, Rohrbach, and Elhoseiny, Buzzega et al.(2020)Buzzega, Boschini, Porrello, Abati, and Calderara, Boschini et al.(2022)Boschini, Bonicelli, Buzzega, Porrello, and Calderara], which stores and reuses data from past tasks; and Architectural [Rusu et al.(2016)Rusu, Rabinowitz, Desjardins, Soyer, Kirkpatrick, Kavukcuoglu, Pascanu, and Hadsell, Madotto et al.(2020)Madotto, Lin, Zhou, Moon, Crook, Liu, Yu, Cho, and Wang], which introduces task-specific parameters to the agent’s architecture. This paper proposes two Rehearsal-based methods, PerpR and ESR, and evaluates their effectiveness through comparative experiments.
@@
Vision-and-Language Navigation. In VLN, agents navigate realistic 3D indoor environments by following natural language instructions. Various datasets like R2R [Anderson et al.(2018b)Anderson, Wu, Teney, Bruce, Johnson, Sünderhauf, Reid, Gould, and Van Den Hengel], R4R [Jain et al.(2019)Jain, Magalhaes, Ku, Vaswani, Ie, and Baldridge], and RxR [Ku et al.(2020)Ku, Anderson, Patel, Ie, and Baldridge] have been developed, focusing on navigation based on initial instructions that fully describe the path. Other tasks, such as CVDN [Thomason et al.(2020)Thomason, Murray, Cakmak, and Zettlemoyer] and HANNA [Nguyen and Daumé III(2019)], involve agents obtaining navigation information through dialogue with an oracle or other agents. Additionally, datasets like Touchdown [Chen et al.(2019)Chen, Suhr, Misra, Snavely, and Artzi] and StreetLearn [Mirowski et al.(2019)Mirowski, Banki-Horvath, Anderson, Teplyashin, Hermann, Malinowski, Grimes, Simonyan, Kavukcuoglu, Zisserman, et al.] have expanded VLN research to outdoor environments. IVLN [Krantz et al.(2023)Krantz, Banerjee, Zhu, Corso, Anderson, Lee, and Thomason] present a paradigm where agents maintain memory across episodes to navigate a sequence of episodes. In the field of VLN, research focuses on improving agent architecture [Hong et al.(2020)Hong, Wu, Qi, Rodriguez-Opazo, and Gould, Chen et al.(2021)Chen, Guhur, Schmid, and Laptev, Fried et al.(2018)Fried, Hu, Cirik, Rohrbach, Andreas, Morency, Berg-Kirkpatrick, Saenko, Klein, and Darrell, An et al.(2024)An, Wang, Wang, Wang, Huang, He, and Wang], enhancing visual-text alignment through auxiliary losses [Ma et al.(2019)Ma, Lu, Wu, AlRegib, Kira, Socher, and Xiong, Wang et al.(2019)Wang, Huang, Celikyilmaz, Gao, Shen, Wang, Wang, and Zhang, Zhu et al.(2020)Zhu, Zhu, Chang, and Liang], using data augmentation [Tan et al.(2019)Tan, Yu, and Bansal, Wang et al.(2022)Wang, Montgomery, Orbay, Birodkar, Faust, Gur, Jaques, Waters, Baldridge, and Anderson, Li et al.(2022)Li, Tan, and Bansal, Kamath et al.(2023)Kamath, Anderson, Wang, Koh, Ku, Waters, Yang, Baldridge, and Parekh], and extensive pre-training [Hao et al.(2020)Hao, Li, Li, Carin, and Gao, Chen et al.(2021)Chen, Guhur, Schmid, and Laptev, Chen et al.(2022)Chen, Guhur, Tapaswi, Schmid, and Laptev, An et al.(2023)An, Qi, Li, Huang, Wang, Tan, and Shao] to enhance generalization. In this paper, we adopted VLN-BERT [Hong et al.(2020)Hong, Wu, Qi, Rodriguez-Opazo, and Gould] and HAMT [Chen et al.(2021)Chen, Guhur, Schmid, and Laptev] as the backbone architectures. Existing VLN datasets and methods lack considerations for CL, highlighting the need for scalable strategies that enable agents to learn efficiently to diverse, changing environments without costly retraining models from scratch.
@@ 3 Continual Vision-and-Lauguage Navigation
3.1 Formulation of VLN and CVLN
In VLN, agents are trained on fixed data and evaluated on episodes from unseen scenes. Given a dataset $D$ of episodes with navigation instructions $I$ and ground-truth trajectories $A^{*}=\{a_{1}^{*},\ldots,a_{N}^{*}\}$, the agent $\pi_{\theta}$ sequentially predicts actions based on visual observations $V=\{v_{1},\ldots,v_{N}\}$ and instructions $I$. The objective minimizes the navigation imitation loss $\ell$, typically cross-entropy:
@@ 3.1 Formulation of VLN and CVLN
CVLN extends this by introducing continual learning across $S$ scene domains. In each domain $s$, episodes are drawn from distribution $D_{s}$, and the agent must learn new domains without forgetting previous ones. The learning objective becomes:
[…]
This continual learning setting better reflects real-world scenarios where agents encounter and must adapt to new environments over time.
@@
3.2 I-CVLN and D-CVLN Benchmarks
We design two CVLN benchmarks: Initial-Instruction based CVLN (I-CVLN) and Dialogue-based CVLN (D-CVLN), constructed from existing VLN datasets. In I-CVLN, agents receive an initial instruction describing the full navigation path. Using the R2R [Anderson et al.(2018b)Anderson, Wu, Teney, Bruce, Johnson, Sünderhauf, Reid, Gould, and Van Den Hengel] and RxR [Ku et al.(2020)Ku, Anderson, Patel, Ie, and Baldridge] datasets, we construct the benchmark with English-only episodes. The training split from these datasets provides 15,700 episodes, while validation uses 1,563 unseen episodes from the same scenes (val-seen split). Episodes are grouped into 20 scene domains, each containing over 600 training episodes. In D-CVLN, agents navigate based on dialogue history. This benchmark is built using the CVDN [Thomason et al.(2020)Thomason, Murray, Cakmak, and Zettlemoyer] dataset, which includes dialogues between a follower and an oracle paired with corresponding trajectories. Due to limited episodes per scene, four scenes are grouped into each scene domain, resulting in 11 domains. The training split includes 3,737 episodes, and validation uses 382 unseen episodes from the same domains. In both benchmarks, scene domains are learned sequentially, aligning with the continual learning setting.
3.3 Evaluation of CVLN
CVLN evaluation measures an agent’s ability to retain prior knowledge while adapting to new scene domains. We use the Average Metric, defined as:
@@ 3.3 Evaluation of CVLN
where $R_{S,i}$ is the agent’s performance on the $i$th domain after learning the $S$th domain. For I-CVLN, $R_{S,i}$ uses standard VLN metrics [Anderson et al.(2018a)Anderson, Chang, Chaplot, Dosovitskiy, Gupta, Koltun, Kosecka, Malik, Mottaghi, Savva, et al.]: Success Rate (SR), Success weighted by inverse Path Length (SPL), and Navigation Error (NE). For D-CVLN, we use Goal Progress (GP) in meters from CVDN [Thomason et al.(2020)Thomason, Murray, Cakmak, and Zettlemoyer].
@@ 4 Methods
Rehearsal-based CL approaches mitigate catastrophic forgetting by using a fixed-size replay memory $M$ to store and revisit episodes from previous domains. After training each domain, $M$ is updated by deleting an equal number of old episodes and adding new ones from the latest domain. Effective replay memory construction and usage are crucial for maintaining performance across domains. In this context, we propose Perplexity Replay (PerpR) and Episodic Self-Replay (ESR) as baseline methods for the CVLN agent. Figure 3 provides an overview of PerpR and ESR.
Figure 3: Overview of Perplexity Replay (PerpR) and Episodic Self-Replay (ESR) for CVLN: (a) PerpR prioritizes challenging episodes with high Action Perplexity (AP), and (b) ESR enables self-replay using past behaviors.